OpenStudy (anonymous):

In triangle ABC, a = 3, b = 5, and c = 7. Find the approximate value of angle A. 22° 38° 142° 158°

OpenStudy (johnweldon1993):

We can use the law of cosine here we know \[\large a^2 = b^2 + c^2 - 2bc \times cos(A)\] So since we want to solve for A we need to rearrange this into \[\large A = \arccos(\frac{a^2 - b^2 - c^2}{-2bc})\] So just plug in everything you have and solve

OpenStudy (anonymous):

what is arccos

OpenStudy (johnweldon1993):

the inverse cosine function also known as \(\large cos^{-1}\)

OpenStudy (anonymous):

okay and i just got .92

OpenStudy (anonymous):

btw i dont have an actual calculator

OpenStudy (johnweldon1993):

The 0.92 is what you get when you do everything in the parenthesis...now you just need the inverse cosine of that Which, since you dont have a calculator...comes out to 21.786

OpenStudy (johnweldon1993):

So that would lead to your "approximate" answer to be 22 degrees